A right triangle has an angle of 8° and the opposite side is 246 cm. What is the length of the hypotenuse?
1
Expert's answer
2013-05-02T07:54:15-0400
Let's denote the first leg as a, so a=246 cm. At first, let's find the length of the second leg b:
b=246*cotan(8°)=1757 cm
Now we can find hypotenuse c. Due to Pythagorean theorem:
c^2=a^2+b^2
Substituting our values we obtain:
c^2=1757^2+246^2=3147565
Therefore c=sqrt(3147565)=1774 cm